OpenStudy (albert0898):

111. The fastest jet airplane can fly at a top speed of about 3,200 feet per second. Approximately how many miles can the jet travel in 15 hours, if it is traveling at top speed? (Note: There are 5,280 feet in one mile.) 112. If k is a positive integer, then for every value of k, the sum of the k smallest distinct odd positive integers is equal to which of the following?

OpenStudy (abb0t):

First, convert feet to miles, then seconds to hours. \(\large\sf \frac{3200 ~ft}{second} \times (\frac{3600~seconds}{hour}) \times \frac{1~mile}{5280 ~ft}\) Your answer should be in the thousands. Just an FYI.